Multi-Variable Calculus Assignment

due Friday, March 23, 2001

Exercises

1. Find the point on the curve

x^{2} − y + z^{2} = 0
xyz = 10

that is nearest to the origin.

2. Find the curvature vector and the curvature of the twisted cubic

t ⟼ (3t, 3t^{2}, 2t^{3})

at the point of that curve corresponding to the parameter value 1.